# Spatially Resolved Kinematics of Post-starburst Galaxies

### v1

@@ 2.3 Other sources of data
We obtain the data for a different sample of 260 early-type galaxies from the ATLAS3D Project (Cappellari et al., 2011). This project obtained its data from the SAURON integral-field unit on the William Herschel Telescope (WHT), combined with deep imaging done with MegaCam on the Canada-France-Hawaii Telescope (CFHT).
We use data sets provided by multiple papers under the umbrella of the ATLAS3D project to obtain parameters that enable us to make comparisons for angular momentum in the context of luminosity-weighted (LW) age, mass-weighted (MW) age, and stellar mass. We acquire ellipticity (Epsilon_e) and angular momentum (LambdaR_e) at effective radius from Emsellem et al. (2011). Emsellem et al. (2011) calculated the $\lambda_{Re}$ using the expression from Emsellem et al. (2007), which is given in Eq 1.
@@
Galaxy - Name of the galaxy (Emsellem et al., 2011)
Lambda_obs - Specific angular momentum as calculated by Emsellem et al. (2011)
Epsilon - Ellipticity (Epsilon_e) from (Emsellem et al., 2011)
FWHM_PSF arcsec Full-width half maximum (dispersion)333used an average value of 1.25” (Scott et al., 2013)
Sersic_n - Sérsic index (n) (Krajnović et al., 2013) of the galactic disk
Semi_Maj_axis kpc Semi-major axis calculated using logRem̂aj and D (Cappellari et al., 2013a) of the galaxy
Lambda_corr - Lambda_obs corrected for beam-smearing (Graham et al., 2018)
Stellar_mass M Stellar mass calculated using logLum (Cappellari et al., 2013a)
and logML_star (Cappellari et al., 2013b) as described in §2.3
LW_age Gyr Light-weighted age (Age_SSP) (McDermid et al., 2015) of the galaxy
MW_age Gyr Mass-weighted age (Age_SHF) (McDermid et al., 2015) of the galaxy
Table 2: ETG properties for the ATLAS3D comparison sample. Full table available online.
@@ 3.1 Angular momentum calculation and correction
Figure 3: Angular momentum at half-light radius, $\lambda_{Re}$ vs. ellipticity at half-light radius, $\epsilon_{Re}$, for 92 PSBs from the MaNGA Survey and 260 ETGs from the ATLAS-3D Project, with point sizes scaled by log stellar mass. The blue and red contour lines represent the 1577 SFs and 964 ETGs from MaNGA respectively. The black dashed line at $\lambda_{Re}=0.1$ separates the slow and fast rotators according to the classical definition, while the black solid line given by $0.31\times\sqrt{\epsilon_{Re}}$ (Emsellem et al., 2011) performs a similar function as the black dashed line, but takes into account the ellipticity of the galaxies in the classification of slow- and fast rotators. We find that a higher fraction of ETGs are slow rotators than the PSB and SF samples, with the SF samples having significantly higher average $\lambda_{Re}$.
We calculate the specific angular momentum for each galaxy using the expression from Emsellem et al. (2007) as:
@@
We calculate the angular momentum (given in column 3 of Table 1) of the 92 PSBs using Eq 1 and the lambdaR_e_calc function defined by Graham et al. (2018) by passing the stellar velocity, stellar velocity dispersion and flux in each pixel (§2). We also provide effective radius, ellipticity, position angle, FWHM_psf, and Sérsic index (all of which are described in Table 1) from the NSA catalog, in order to determine the 1 $R_{e}$ ellipse in which to calculate $\lambda$. The stellar velocity dispersion has been corrected for the instrumental dispersion (stellar_sigmacorr) using the given correction provided by Westfall et al. (2019):
@@
According to Graham et al. (2018), the finite point-spread function (PSF) results in the smearing of the velocity field. To correct for this effect, they provide an optional analytic correction as part of their lambdaR_e_calc function. This correction can be applied to regular rotators whose semi-major axis is larger than the PSF dispersion (sig_PSF) and Sérsic index is in the range $0.5-6.5$. In order to apply this correction to all the galaxies in our sample, we replace the Sérsic indices with 0.5001 if $n<0.5$ and with 6.4999 if $n>6.5$. Thus, we obtain the beam-corrected angular momentum for the 92 PSBs. All the parameters used and the angular momentum calculated for the PSB are presented in Table 1.
@@ 3.2 Identifying slow and fast rotators
We plot the beam-corrected specific angular momentum at half-light radius for 92 MaNGA PSBs and 260 ETGs from ATLAS3D against the ellipticity at half-light radius in Figure 3. The points are scaled in size and color by stellar mass (acquired as described in §2.1 §2.3 for PSBs and ETGs respectively), with the corresponding mass given in the legend. The blue and red contours depict the 1577 star-forming and 964 early-type MaNGA galaxies respectively whose data is obtained from Graham et al. (2018). We find that a higher fraction of ETGs are slow rotators than the PSB and SF samples, with the SF samples having significantly higher average $\lambda_{Re}$.
@@
As shown in Figure 3, we show two criteria to separate the slow and fast rotators: the classical $\lambda_{Re}=0.1$ (black dashed line) boundary condition and the curve given by $0.31\times\sqrt{\epsilon_{Re}}$ (black dotted line) (Emsellem et al., 2011). Emsellem et al. (2011) found that the specific angular momentum generally increases with the ellipticity, even for non-regular rotators, so they developed a relationship between $\lambda_{R}-\epsilon$ to distinguish slow and fast rotators. This relationship also accounts for the increasing galactic anisotropy with ellipticity. Thus, $\lambda_{Re}=0.31\times\sqrt{\epsilon_{Re}}$ is a better measure of whether a galaxy is a slow or fast rotator. Using this as the preferred limiting condition, we find that 5 out of 92 ($\approx$5%) PSBs, 36 out of 260 ($\approx$14%) ETGs from ATLAS3D, 64 out of 1577 ($\approx$4%) MaNGA SFs, and 213 out of 964 ($\approx$22%) MaNGA ETGs are slow rotators. We also find that $84\%$ of SFs have higher angular momentum than the maximum $\lambda_{Re}$ of the PSB sample. Thus, the ETGs generally have lower angular momentum as compared to the PSBs, which tend to rotate slower than the SFs.
@@
The scaling and coloring of points based on stellar mass shows that galaxies, especially ETGs, with higher stellar mass prefer to rotate slower as compared to their counterparts. We explore trends in angular momentum with stellar mass in Figure 4. In this Figure, we can see a distinct population of high stellar mass slow rotators in the ETG sample, that is not present for the PSBs and SF galaxies. To quantify this observation, we consider the fraction of slow rotators in the top 30th percentile of stellar mass for each galaxy type. $42\%$ of MaNGA SFs, $40\%$ of MaNGA PSBs, $58\%$ of MaNGA ETGs and $58\%$ of ATLAS3D ETGs that are slow rotators, are in the top 30th percentile of stellar mass. While the slow rotator SF and PSB galaxies are slightly over-represented in the highest stellar mass galaxies, the slow rotator ETGs are significantly more biased towards the higher mass galaxies. We find similar levels of over-representation when we consider the top 10th or 20th percentiles of stellar mass.
@@
We consider the fraction of slow rotators in the PSBs within the different parent samples. Of the 92 PSBs, there are 25 E+As, 19 SPOGs, and 77 PCAs (with significant overlap). Two of the 5 slow rotators are in multiple classifications, and we find 1 E+A, 1 SPOG, and 5 PCA slow rotators. Within the small number uncertainties, the fraction of slow rotators is consistent between each parent PSB sample.
@@
Figure 4: Angular momentum vs. stellar mass for 92 PSBs (black scatter), 1577 SFs (blue contours) and 964 ETGs (red contours) from MaNGA, and 260 ATLAS3D ETGs (red scatter). We find that galaxies, especially the early-types, with higher stellar mass tend to rotate slowly as compared to galaxies with lower stellar mass. While most of the slow rotator ETGs are amongst the most massive ETGs, a similar effect is not seen for the PSB and SF samples. We explore this effect quantitatively in the text. This implies that the high stellar mass slow rotator ETGs have experienced further growth and loss of angular momentum since they may have passed through the PSB phase.
@@ 4.1 Evolution of angular momentum with age
Figure 5: Left: Angular momentum at half-light radius, $\lambda_{Re}$, vs. mass-weighted age for 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for PSBs. We observe a clear distinction in age between the SF and ETG samples, with PSB galaxies having the youngest MW ages. A significant anti-correlation between MW age and $\lambda_{Re}$ is seen for the SF and ETG samples, though not within the PSB sample. At mass-weighted ages $\sim$5 Gyr, where the oldest PSBs have similar mass-weighted ages to the youngest ETGs, the PSB and ETG distributions in angular momentum overlap. For older ETGs, we see that a trend towards lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. Right: Angular momentum at half-light radius, $\lambda_{Re}$, vs. light-weighted age 92 PSBs (black points), 1352 SFs (blue contours), and 903 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. Similar to Figure 5a, we observe a clear distinction between the SF and ETG samples, though now in the case of LW age, the PSB ages are typically between the SF and ETG samples. For the LW ages, we observe a significant trend within the ETG samples, but a weaker trend in the SF sample and no significant trend in the PSB sample. The fact that the angular momentum is more closely coupled with the mass-weighted age implies that it is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age.
@@
To further our understanding of the evolution of angular momentum across the three classes, we study its evolution with the age of the stellar populations. To do so, we plot the angular momentum against mass- and luminosity-weighted age (as shown in Figure 5. As discussed in §2.2, we obtain the respective ages for MaNGA SFs, PSBs and ETGs from SDSS Release 17 pyPipe3D pipeline (v3_1_1), which is a result of the analysis of stellar populations and ionized gas as presented in Sánchez et al. (2016a, b, 2018).
@@
We observe that angular momentum decreases with increasing age across all three samples for MW and LW age, with the early types having older MW and LW ages and more slow rotators than PSBs and SFs. At mass-weighted ages $\sim$5 Gyr, where the oldest PSBs have similar mass-weighted ages to the youngest ETGs, the PSB and ETG distributions in angular momentum overlap. For older ETGs, we see that a trend towards lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. To quantify this, we compute the Spearman correlation (presented in Table 3) for all three classes of galaxies for MW and LW age. We find significant correlations of angular momentum with mass-weighted age for the SF and ETG samples, though not within the PSB sample. For the light-weighted ages, we see a weaker correlation for the SF sample. The fact that the angular momentum is more closely coupled with the mass-weighted age implies that it is more closely coupled to the longer-scale formation history of the galaxy, instead of the more recent history traced by the light-weighted age.
@@
We also test whether we see evolution within the PSB phase by testing the correlation between angular momentum and age since the starburst ended using the stellar population fits by French et al. (2018). Consistent with the lack of correlation in MW or LW age, we observe no significant trend within the PSB in angular momentum vs. starburst age.
@@
Table 3: Spearman result for MW and LW ages for MaNGA SFs, PSBs, and ETGs
@@
SFs ($\rho$ = -0.22, p $\approx 10^{-16}$) ($\rho$ = -0.06, p = 0.04)
PSBs ($\rho$ = -0.11, p = 0.31) ($\rho$ = -0.16, p = 0.13)
ETGs ($\rho$ = -0.33, p $\approx 10^{-24}$) ($\rho$ = -0.29, p $\approx 10^{-19}$)

### v2

@@ 2.3 Other Sources of Data
We obtain the data for a different sample of 260 ETGs from the ATLAS3D Project (Cappellari et al., 2011). This project obtained its data from the SAURON IFU on the William Herschel Telescope (WHT), combined with deep imaging done with MegaCam on the Canada-France-Hawaii Telescope.
We use data sets provided by multiple papers under the umbrella of the ATLAS3D project to obtain parameters that enable us to make comparisons for angular momentum in the context of LW age, MW age, and stellar mass. We acquire ellipticity (Epsilon_e) and angular momentum (LambdaR_e) at effective radius from Emsellem et al. (2011). Emsellem et al. (2011) calculated the $\lambda_{Re}$ using the expression from Emsellem et al. (2007), which is given in Equation (1).
@@
Table 2: ETG properties for the ATLAS3D Comparison Sample [^6] This table is available in its entirety in machine-readable form in the online article.
@@
Galaxy Name of the galaxy (Emsellem et al., 2011)
Lambda_obs Specific angular momentum as calculated by Emsellem et al. (2011)
Epsilon Ellipticity (Epsilon_e) from (Emsellem et al., 2011)
[…]
Sérsic_n Sérsic index (n) of the galactic disk (Krajnović et al., 2013)
Semi_Maj_axis kpc Semimajor axis calculated using logRem̂aj and D of the galaxy (Cappellari et al., 2013a)
Lambda_corr Lambda_obs corrected for beam-smearing (Graham et al., 2018)
Stellar_mass M Stellar mass calculated using logLum (Cappellari et al., 2013a) and
logML_star (Cappellari et al., 2013b) as described in Section 2.3
LW_age Gyr Luminosity-weighted age (Age_SSP) of the galaxy (McDermid et al., 2015)
MW_age Gyr Mass-weighted age (Age_SHF) of the galaxy (McDermid et al., 2015)
@@ 3.1 Angular Momentum Calculation and Correction
Figure 3: Angular momentum at half-light radius, $\lambda_{Re}$ vs. ellipticity at half-light radius, $\epsilon_{Re}$, for 89 PSBs from the MaNGA Survey and 260 ETGs from the ATLAS3D Project, with point sizes scaled by log stellar mass. The blue and red contour lines represent the 1592 SFGs and 916 ETGs from MaNGA, respectively. The black dashed line at $\lambda_{Re}=0.1$ separates the slow and fast rotators according to the classical definition, while the black solid line given by $0.31\times\sqrt{\epsilon_{Re}}$ (Emsellem et al., 2011) performs a similar function as the black dashed line, but takes into account the ellipticity of the galaxies in the classification of slow and fast rotators. We find that a higher fraction of ETGs are slow rotators than the PSB and SFG samples, with the SFG sample having significantly higher average $\lambda_{Re}$.
We calculate the specific angular momentum for each galaxy using the expression from Emsellem et al. (2007) as
@@
We calculate the angular momentum (given in column (3) of Table 1) of the 89 PSBs, 1592 SFGs and 916 ETGs using Equation (1) and the lambdaR_e_calc function defined by Graham et al. (2018) by passing the stellar velocity, stellar velocity dispersion, and flux in each pixel (Section 2). We also provide effective radius, ellipticity, position angle, full width at half-maximum (FWHM), and Sérsic index (all of which are described in Table 1) from the NSA catalog, in order to determine the 1 $R_{e}$ ellipse in which to calculate $\lambda$. The stellar velocity dispersion has been corrected for the instrumental dispersion (stellar_sigmacorr) using the given correction provided by Westfall et al. (2019):
@@
According to Graham et al. (2018), the finite point-spread function (PSF) results in the smearing of the velocity field. To correct for this effect, they provide an optional analytic correction as part of their lambdaR_e_calc function. This correction can be applied to regular rotators whose semimajor axis is larger than the PSF dispersion and Sérsic index is in the range $0.5-6.5$. Thus, we obtain the beam-corrected angular momentum for the MaNGA galaxies. All the parameters used and the angular momentum calculated for the MaNGA SFGs, PSBs and ETGs are presented in Table 1.
@@ 3.2 Identifying Slow and Fast Rotators
We plot the beam-corrected specific angular momentum at half-light radius for 89 MaNGA PSBs and 260 ETGs from ATLAS3D against the ellipticity at half-light radius in Figure 3. The points are scaled in size and color by stellar mass (acquired as described in Section 2.1 and Section 2.3 for PSBs and ETGs, respectively), with the corresponding mass given in the legend. The blue and red contours depict the 1592 SFG and 916 ETG MaNGA galaxies, respectively. We find that a higher fraction of ETGs are slow rotators than the PSB and SFG samples, with the SFG samples having significantly higher average $\lambda_{Re}$.
@@
As shown in Figure 3, we show two criteria to separate the slow and fast rotators: the classical $\lambda_{Re}=0.1$ (black dashed line) boundary condition and the curve given by $0.31\times\sqrt{\epsilon_{Re}}$ (black dotted line) (Emsellem et al., 2011). Emsellem et al. (2011) found that the specific angular momentum generally increases with the ellipticity, even for non-regular rotators, so they developed a relationship between $\lambda_{Re}-\epsilon_{Re}$ to distinguish slow and fast rotators. This relationship also accounts for the increasing galactic anisotropy with ellipticity. Thus, $\lambda_{Re}=0.31\times\sqrt{\epsilon_{Re}}$ is a better measure of whether a galaxy is a slow or fast rotator. Using this as the preferred limiting condition, we find that 5 out of 89 ($\sim$6%) PSBs, 36 out of 260 ($\sim$14%) ETGs from ATLAS3D, 56 out of 1592 ($\sim$3.5%) MaNGA SFGs, and 196 out of 916 ($\sim$19.5%) MaNGA ETGs are slow rotators. We also find that $\sim 42\%$ of MaNGA SFGs, $\sim 25\%$ of MaNGA PSBs, $\sim 18\%$ of MaNGA ETGs and $\sim 21\%$ of ATLAS3D ETGs have higher angular momentum than $\lambda_{Re}=0.6$. Furthermore, within the MaNGA sample, $\sim 64\%$ SFGs have higher momentum than the median $\lambda_{Re}\approx 0.5$ of the fast-rotator PSBs, and $\sim 74\%$ SFGs and $\sim 62\%$ PSBs rotate faster than the median $\lambda_{Re}\approx 0.44$ of the fast-rotator ETGs. The median $\lambda_{Re}\approx 0.57$ of the SFGs is higher than the median $\lambda_{Re}$ of the fast-rotator PSBs and ETGs, respectively. Thus, the ETGs generally have lower angular momentum as compared to the PSBs, which tend to rotate slower than the SFGs.
@@
The scaling and coloring of points based on stellar mass shows that galaxies, especially ETGs, with higher stellar mass prefer to rotate slower as compared to their counterparts. We explore trends in angular momentum with stellar mass in Figure 4. In this figure, we can see a distinct population of high-stellar-mass slow rotators in the ETG sample that is not present for the PSBs and SFGs. To quantify this observation, we consider the fraction of slow rotators in the top 30th percentile of stellar mass for each galaxy type: $\sim 32\%$ of MaNGA SFGs, $40\%$ of MaNGA PSBs, $\sim 58\%$ of MaNGA ETGs and $\sim 58\%$ of ATLAS3D ETGs. While the slow-rotator SFG and PSB galaxies are slightly overrepresented in the highest-stellar-mass galaxies, the slow-rotator ETGs are significantly more biased towards the higher-mass end. We find similar levels of overrepresentation when we consider the top 10th or 20th percentiles of stellar mass.
@@
We consider the fraction of slow rotators in the PSBs within the different parent samples. Of the 89 PSBs, there are 24 E+As, 19 SPOGs, and 74 PCAs (with significant overlap). Two of the five slow rotators are in multiple classifications, and we find one E+A, one SPOG, and five PCA slow rotators. Within the small number uncertainties, the fraction of slow rotators is consistent between each parent PSB sample.
@@
Figure 4: Angular momentum vs. stellar mass for 89 PSBs (black scatter), 1592 SFGs (blue contours) and 916 ETGs (red contours) from MaNGA, and 260 ATLAS3D ETGs (red scatter). The black line represents the running average for the PSBs. We find that galaxies, especially the early types, with higher stellar mass tend to rotate slowly as compared to galaxies with lower stellar mass. While most of the slow-rotator ETGs are among the most massive ETGs, a similar effect is not seen for the PSB and SFG samples. We explore this effect quantitatively in the text. This implies that the high-stellar-mass slow-rotator ETGs have experienced further growth and loss of angular momentum since they may have passed through the PSB phase.
@@ 4.1 Evolution of Angular Momentum with Age
Figure 5: Left: angular momentum at half-light radius, $\lambda_{Re}$, vs. mass-weighted (MW) age for 89 PSBs (black points), 1586 SFGs (blue contours), and 914 ETGs (red contours) from the MaNGA survey. The black line represents the running average for PSBs. We observe a clear distinction in age between the SFG and ETG samples, with the median MW age of the PSB sample ($\sim$3.4 Gyr) younger than the median MW ages for the SFG ($\sim$4.6 Gyr) and ETG ($\sim$6.9 Gyr) samples. A significant anticorrelation between MW age and $\lambda_{Re}$ is seen for the SFG and ETG samples, though not within the PSB sample. At MW ages $\sim$5 Gyr, where we observe an overlap between the older PSBs and the younger ETGs, the PSB and ETG distributions in angular momentum are quite similar. For older ETGs, we see that a trend towards lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. Right: angular momentum at half-light radius, $\lambda_{Re}$, vs. luminosity-weighted (LW) age for 89 PSBs (black points), 1586 SFGs (blue contours), and 914 ETGs (red contours) from the MaNGA survey. The black line represents the running average for the PSBs. Similar to Figure 5(a), we observe a clear distinction between the SFG and ETG samples, though now in the case of LW age, the PSB ages are typically between the SFG and ETG samples. Quantitatively, the median LW age of the PSB sample ($\sim$1.45 Gyr) is older than the median LW age for the SFGs ($\sim$0.9 Gyr) and younger than the median LW age of the ETGs ($\sim$4.9 Gyr). For the LW ages, we observe a significant trend within the ETG samples, but no significant trend in the SFG and the PSB sample. The fact that the angular momentum is more closely coupled with the MW and LW age for the SFGs and ETGs as compared to the PSBs implies that the kinematic evolution is more closely coupled to the longer-scale formation history and evolution of the galaxy in its star-forming and early-type phases, instead of the time it spends in the post-starburst phase.
@@
To further our understanding of the evolution of angular momentum across the three classes, we study its evolution with the age of the stellar populations. To do so, we plot the angular momentum against MW and LW ages (as shown in Figure 5). As discussed in Section 2.2, we obtain the respective ages for MaNGA SFGs, PSBs and ETGs from the SDSS DR17 pyPipe3D pipeline (v3_1_1), which is a result of the analysis of stellar populations and ionized gas as presented in Sánchez et al. (2016a, b, 2018).
@@
We observe that angular momentum decreases with increasing age across all three samples for MW and LW age, with the early types having older MW and LW ages and more slow rotators than PSBs and SFGs. While the PSBs are on average younger than both the SFGs and ETGs based on their median MW age, their median LW age seems to be older than the average SFG. The younger MW age of the PSBs as compared to the SFGs and ETGs can be attributed to the significant recent burst populations in these galaxies, while the recent quenching in the PSBs leads to their older LW ages as compared to the SFGs. At MW ages $\sim$5 Gyr, where the older PSBs have similar MW ages to the younger ETGs, the PSB and ETG distributions in angular momentum overlap. For older ETGs, we see that a trend toward lower angular momentum continues. This implies that even after the PSBs evolve into ETGs, they do not stop shedding off their angular momentum. To quantify this, we compute the Spearman correlation (presented in Table 3) for all three classes of galaxies for MW and LW age. We find significant anticorrelations of angular momentum with MW age for the SFG and ETG samples, though not within the PSB sample. For the LW age, we do not observe a significant anticorrelation for the PSB sample as well. The fact that the angular momentum is more closely coupled with the MW and LW ages for the SFGs and ETGs as compared to the PSBs implies that the kinematic evolution is more closely coupled to the longer-scale formation history and evolution of the galaxy in its star-forming and early-type phases, instead of the time it spends in the post-starburst phase.
@@
We also test whether we see evolution within the PSB phase by testing the correlation between angular momentum and age since the starburst ended using the stellar population fits by French et al. (2018). Consistent with the lack of correlation in MW or LW ages, we observe no significant trend within the PSBs in angular momentum versus starburst age.
@@
Table 3: Spearman Result for MW and LW Ages for MaNGA SFGs, PSBs, and ETGs
@@
SFGs ($\rho\approx$0.12, p $\approx$ $10^{-6}$) ($\rho\approx$ $0.11$, p $\approx 10^{-5}$)
PSBs ($\rho\approx$0.13, p $\approx$ 0.22) ($\rho\approx$0.15, p $\approx$ 0.16)
ETGs ($\rho\approx$0.30, p $\approx 10^{-20}$) ($\rho\approx$0.28, p $\approx 10^{-18}$)